OpenStudy (anonymous):

ANOTHER. :D x+3y=-5 2x+y=5

OpenStudy (anonymous):

x = -5 -3y 2( -5 - 3y ) +y = 5 -10 -6y + y = 5 -5y = 5 + 10 -5y = 15 y = -3 x + 3y = -5 x + 3(-3) = -5 x -9 = -5 x = 4

OpenStudy (anonymous):

It's so nice to solve other's problems. XDD

OpenStudy (anonymous):

I know haha. Too bad I can't solve my own. :'(

OpenStudy (anonymous):

Thanks for the help Josephniel14 ;)

OpenStudy (anonymous):

My pleasure. :))